a runaway train has a mass of 12,000 kg and travels at the speed of 5.0 ms down a track. Complete the time required for a force of 1000 N force to flow the car down to1.0

48 s

60 s

72 s

120 s
Answer #1

By impulse momentum relation,

we know that impulse given = change in momentum

-F*t = 0-mv

1000*t = 12000*5

t = 12000*5/1000

= 60 s answer